# Evaluation of machine-learning models to measure individualized treatment effects from randomized clinical trial data with time-to-event outcomes

### v1

@@ 3.4.1 Breast cancer study on taxane chemotherapy
We apply the models to a meta-analysis of neoadjuvant trials that included $614$ breast cancer patients treated by anthracycline-based chemotherapy with ($n=507$) or without ($n=107)$ taxane. The dataset is composed of $p=1,689$ genes. This cohort is available in the biospear(Ternès et al., 2018) R package that is publicly available.
@@ 3.4.2 Breast cancer cohort with adjuvant trastuzumab
The models are also compared on data from an RCT funded by the National Cancer Institute and the National Institute of Health. This randomized phase III clinical trial evaluated the effect of adding trastuzumab to adjuvant chemotherapy to disease-free survival in early breast cancer patients in Her2+ breast cancer.
$n=1,574$ patients were included in the trial: $n=795$ patients received chemotherapy, while $n=779$ others received chemotherapy and trastuzumab. For each patient, the expression of $462$ genes is available in addition to standard clinicopathological variables such
@@ 3.5 Evaluation Measures
The treatment effect can be defined as the effect of treatment $\tau$ on the outcome variable $Y$. Since $Y$ is the observed outcome, we have:
In an RCT, assuming that $Y(1)$ and $Y(0)$ are independent from $\tau|X$, the treatment effect is defined as:
with $X$ the characteristics of the patient.
[…]
@@ 3.5.1 Observed and predicted treatment benefit
We apply $2$ matching procedures, either using variables or predicted benefit. First, we use a predicted score for each individual. The $2$ groups of patients (control arm and treatment arm) are ranked based on the value of the predicted benefit. The data points are matched by aligning the sorted lists based on their ranks. Secondly, we apply a matching procedure based on patients’ characteristics and the Mahalanobis distance. It is the distance between $2n$ dimensional points scaled by the statistical variation in each component of the point.
The predicted benefit $\hat{\theta}$ is defined as the difference between the individual survival probability at a given timepoint with the treatment option being tested minus the probability with the control for the same patient. Given that matching is not perfect, the predicted benefits are averaged for the pair of patients previously matched, and predicted benefit is defined as the average of the predicted benefits within each matched patient pair. We have :
$\hat{\theta}_{u,v}(t)=\frac{[S_{u}(t|\tau=1,X_{u})-S_{u}(t|\tau=0,X_{u})]}{2}+\frac{[S_{v}(t|\tau=1,X_{v})-S_{v}(t|\tau=0,,X_{v})]}{2},$ (11)
@@
The first term of the equation $\mathbf{1}_{Statut_{i}}\times\mathbf{1}_{\tilde{T}_{u}<\tilde{T}_{v}}$ equals $1$ if the patient $i$ has a shorter time-to-event compared to patient $v$, and 00 otherwise.The reasoning is the same for the second term of the equation (for patient $v$). When the shortest survival time is censored, the treatment effect is not observed for the patient pair and equals 00.
Thus, for each pair of patients, the observed treatment benefit can take one of the following values: $1$ for observed treatment benefit, 00 for no observed treatment effect, and $-1$ for deleterious treatment effect.
@@ 3.5.2 Discrimination measures
The value of the C-statistic lies between 00 and $1$, with $0.5$ equivalent to a random prediction and $1$ corresponding to a perfect ability to rank. We compute the C-statistic for all individuals.
C-for-benefit ($\text{C}_{\text{benefit}}$) was introduced by Klaveren et al. (2017) as a measure of the discrimination capacity of the model. It is an extension of the C-statistic adapted to measure treatment benefit at a horizon time $t$. It represents the probability that, among $2$ randomly chosen matched patient pairs, the pair with greater observed benefit also has a higher mean predicted benefit. It can be written as:
@@ 3.5.3 Calibration measures
Calibration can be assessed by a smoothed calibration curve obtained by a local regression of observed pairwise treatment effect on predicted pairwise treatment effect. The observed benefits are regressed on the predicted benefits using a locally weighted scatterplot smoother (loess). It is a non-parametric regression technique used to create a smoothing function through a scatterplot of data points. It combines multiple local regressions to fit a curve to subsets of the data, allowing it to capture complex patterns without assuming a specific global model.
[…]
@@ 3.5.4 Root mean squared error
with $\theta_{i}$ the true treatment benefit for individual $i$ at time $t$. $\hat{\theta}_{i}$ is the predicted benefit for individual $i$, that is, the difference in survival probabilities at a pre-specified time point $t$ for one patient with and without treatment:
@@
3.5.5 Bootstrap confidence intervals
Expected survival probabilities are estimated at specific time points (at $t=2$ and $t=5$ years), and bootstrap confidence intervals (CI) are constructed following the approach proposed by Roblin et al. (2024).
With this approach, the data is divided into a training set and a test set. $M=200$ bootstrap sets are sampled with replacement from the training set. The model is trained for each of these $M$ subsets. It allows for the estimation of $M$ survival probabilities at time $t$ for a given patient $i$ from the test set, noted as $\hat{S}_{i}(t)=\{\hat{S}_{i(1)}(t),\dots,\hat{S}_{i(M)}(t)\}$. The percentile method is then used to obtain confidence intervals at level $1-\theta$ for $\hat{S}(t)$ based on the distribution of $\hat{S}_{\text{all}}(t)$:
$\text{CI}_{1-\theta}=\Big{[}q_{\frac{\theta}{2}}\Big{(}\hat{S}_{i}(t)\Big{)},q_{1-\frac{\theta}{2}}\Big{(}\hat{S}_{i}(t)\Big{)}\Big{]}$ (16)
where $q_{\frac{\theta}{2}}$ and $q_{1-\frac{\theta}{2}}$ are percentiles computed using the empirical distribution of the $M$ values. Here, $\theta=5\%$ and the $2.5^{th}$ and $97.5^{th}$ percentiles are computed using the empirical distribution of the $M$ survival probabilities.
We calculate the CI for the survival probabilities of a patient $i$ in the treatment arm ($\hat{S}_{i,trt}(t)$) and the matched patient $j$ in the control arm ($\hat{S}_{j,ctrl}(t)$). The matching procedure is based on patients’ characteristics to ensure consistent matched pairs across all models.
@@ 4.1.1 Data generation process 1
In Table 1, the highest average C-statistic and $\text{C}_{\text{benefit}}$ were obtained with the IF model. It was also the closest to the oracle value. ALASSO obtained the lowest value in terms of $\text{C}_{\text{benefit}}$, which could indicate that the linear ALASSO model has difficulty capturing the treatment by biomarker interactions, whereas IF, due to its specific design for such interactions, demonstrates superior performance in modeling them.
[…]
@@
[…]
@@
We also compared the models at different time points (Supplementary material Table 8), and for different levels of censoring (Supplementary material Table 9, Setting 1). All the results were improved when the censoring rate was lower ($20\%$ instead of $50\%$). Regarding the choice of the horizon time, the oracle values decreased over time, while the models performed similarly between the first and third quartiles of the time range.
The models were also compared using a matching procedure based on patient’s characteristics (Table 9, Setting 2). All the models performed worse compared to using a matching procedure based on predicted treatment benefit. As noted by Klaveren et al. (2017), this is expected as the predicted treatment benefit for $2$ patients within a pair matched by patient characteristics tends to be less similar.
@@ 4.1.2 Data generation process 2
[…]
Table 2 also presents the the results for the model trained on a larger sample size. The results demonstrate only subtle differences across the evaluated metrics compared to the results obtained with a smaller sample size.
@@
[…]
@@
4.2 Results on the two breast cancer data sets
FNN-based methods, IF and ALASSO were applied on $2$ breast cancer clinical trial data sets. All comparison measures were computed on the test set at $2$ and $5$ years. These times were chosen as these are usual times of interest for clinical investigators for early breast cancer.
4.2.1 Individualized treatment effects of taxane chemotherapy
[…]
Table 3: Measures at 2 and 5 years for the meta-analysis the effect of taxane chemotherapy. - Matching on predicted benefit.
@@ 4.2.1 Individualized treatment effects of taxane chemotherapy
[…]
ALASSO $\mathbf{0.733}$ $0.605$ $0.266$ $0.267$ $\mathbf{0.703}$ $\mathbf{0.603}$ $0.240$ $0.203$
CoxCC $0.658$ $\mathbf{0.619}$ $0.07$ $\mathbf{0.076}$ $0.550$ $0.599$ $\mathbf{0.069}$ $\mathbf{0.094}$
CoxTime $0.689$ $0.609$ $0.086$ $0.103$ $0.550$ $0.579$ $0.094$ $0.100$
IF $0.694$ $0.587$ $\mathbf{0.032}$ $0.108$ $0.549$ $0.571$ $0.099$ $0.339$
@@
Two matched patients were sampled from the dataset. Patient 1 is part of the control group, with a survival time of 1.5 years without censoring. Patient 2 is part of the treatment group, followed for 2.5 years before being lost to follow-up. Among the models, ALASSO demonstrates the narrowest CIs, indicating higher precision in its predictions, mimicking a bit the low RMSE from the simulation study, but without having the best c-for-benefit or calibration measures. Across all models, Patient 2 consistently shows a higher predicted mean survival probability than Patient 1, with narrower CI. For example, with the CoxTime model, Patient 1 has a mean survival probability of $0.820$ and a CI of $[0.558,0.953]$, whereas Patient 2 has a mean survival probability of $0.942$ with a CI of $[0.858,0.953]$. Figure 4 illustrates the distribution of predicted survival probabilities, showing a wider spread for Patient 1 compared to Patient 2.
Table 4: Mean expected survival probabilities and associated CI at 2 years for 2 patients from the breast cancer cohort on taxane chemotherapy.
Patient ALASSO CoxCC CoxTime IF
Patient 1 with control $\underset{[0.845,0.848]}{0.846}$ $\underset{[0.668,0.968]}{0.857}$ $\underset{[0.558,0.953]}{0.820}$ $\underset{[0.767,0.778]}{0.772}$
Patient 2 with treatment $\underset{[0.892,0.895]}{0.894}$ $\underset{[0.829,0.989]}{0.939}$ $\underset{[0.858,0.983]}{0.942}$ $\underset{[0.925,0.931]}{0.928}$
Figure 4: Density estimate of expected survival probabilities at 2 years computed with CoxTime for two patients with breast cancer.
4.2.2 Individualized treatment effects of adjuvant trastuzumab
For the second breast cancer data set, we observed in Table 5 that ALASSO obtained the highest $\text{C}_{\text{benefit}}$. This could suggest that there are limited nonlinear and high-order interactions with the treatment in the data set, and that the more flexible machine learning method don’t capture a stronger signal. On average, ALASSO selected 22 biomarker main effects and 16 biomarker-by-treatment interactions.
[…]
Table 5: Measures at 2 and 5 years in the RCT on adjuvant trastuzumab. - Matching on predicted benefit.

### v2

@@ 3.4.1 Breast cancer study on taxane chemotherapy
We apply the models to a meta-analysis of neoadjuvant trials that included $614$ breast cancer patients treated by anthracycline-based chemotherapy with ($n=507$) or without ($n=107)$ taxane. The dataset comprises $p=1,689$ genes. This cohort is available in the biospear (Ternès et al., 2018) R package, which is publicly available.
@@ 3.4.2 Breast cancer cohort with adjuvant trastuzumab
The models are also compared on data from an RCT funded by the National Cancer Institute and the National Institute of Health. This randomized phase III clinical trial evaluated the effect of adding trastuzumab to adjuvant chemotherapy on disease-free survival in early breast cancer patients with HER2+ tumors.
$n=1,574$ patients were included in the trial: $n=795$ patients received chemotherapy, while $n=779$ others received chemotherapy and trastuzumab. For each patient, the expression of $462$ genes is available, in addition to standard clinicopathological variables such as ER status or the number of positive nodes.
@@
3.4.3 Advanced Renal Cell Carcinoma cohort with avelumab
The third cohort analyzed comprises baseline tumor samples from the phase 3 JAVELIN Renal 101 trial, which assessed first-line avelumab plus axitinib versus sunitinib alone for progression-free survival in previously untreated patients with advanced renal cell carcinoma (Motzer et al., 2020). A total of $n=886$ patients were enrolled, with $n=443$ assigned to avelumab plus axitinib and $n=443$ to sunitinib. Consistent with the original study, we utilized the provided pathway scores, where each pathway score for a given sample was computed as the average of the standardized expression values of the genes comprising that pathway. The dataset features $58$ pathway scores alongside $2$ clinical variables: age and sex.
Figure 3: Kaplan–Meier survival curves for the advanced renal cell carcinoma cohort from the phase III JAVELIN Renal 101 trial.
@@ 3.5 Evaluation Measures
The models are compared using metrics adapted to the evaluation of treatment benefit: the Concordance index for benefit ($\text{C}_{\text{benefit}}$), the Integrated Calibration Index for benefit ($\text{ICI}_{\text{benefit}}$), the E50 for benefit ($\text{E50}_{\text{benefit}}$), and the root-mean-squared error (RMSE).
@@ 3.5.1 Observed and predicted treatment benefit
Since the counterfactual response for a given patient cannot be observed, the matched patient serves as a surrogate for this unobserved outcome. To construct such pairs, we use two matching procedures, based either on patients’ baseline covariates or on their predicted treatment benefit. For the first method, we compute a predicted benefit score for each individual. Then, within each arm (control and treatment), patients are ranked according to this score, and pairs are formed by aligning the two ranked lists so that patients with similar predicted benefit are matched. For the second method, we apply a covariate-based matching procedure based on patients’ characteristics and the Mahalanobis distance. For each treated patient $u$, we search among all control patients $v$ and identify the one whose covariate vector $X_{u}$ is closest to $X_{v}$ in terms of Mahalanobis distance. We first compare the two matching strategies on the simulation data, and subsequently retain and apply the predicted-benefit–based matching approach in the real-data analysis.
Importantly, we define the conditional average treatment benefit, $\hat{\theta}$, as an estimate of the CATE at time $t$, computed as the difference between the individual survival probability at a given time point with the treatment option being tested minus the probability with the control for the same patient. Given that matching is not perfect, the predicted benefits are averaged for the pair of patients previously matched, and the predicted benefit is defined as the average of the predicted benefits within each matched patient pair. We have :
$\hat{\theta}_{u,v}(t)$ $=\frac{[S_{u}(t|\tau=1,X_{u})-S_{u}(t|\tau=0,X_{u})]}{2}$ (11)
$+\frac{[S_{v}(t|\tau=1,X_{v})-S_{v}(t|\tau=0,,X_{v})]}{2},$
@@
The first term of the equation $\mathbf{1}_{Statut_{i}}\times\mathbf{1}_{\tilde{T}_{u}<\tilde{T}_{v}}$ equals $1$ if the patient $i$ has a shorter time-to-event compared to patient $v$, and 0 otherwise. The reasoning is the same for the second term of the equation (for patient $v$). When the shortest survival time is censored, the treatment effect is not observed for the patient pair and equals 0.
Thus, for each pair of patients, the observed treatment benefit can take one of the following values: $1$ for observed treatment benefit, 0 for no observed treatment effect, and 1-1 for deleterious treatment effect.
@@ 3.5.2 Discrimination measures
The value of the C-statistic lies between 0 and $1$, with $0.5$ equivalent to a random prediction and $1$ corresponding to a perfect ability to rank. We compute the C-statistic for all individuals.
C-for-benefit ($\text{C}_{\text{benefit}}$) was introduced by Klaveren et al. (2017) as an extension of the C-statistic to measure treatment benefit at a horizon time $t$. It represents the probability that, among $2$ randomly chosen matched patient pairs, the pair with greater observed benefit also has a higher mean predicted benefit. It can be written as:
@@ 3.5.3 Calibration measures
Calibration can be assessed by a smoothed calibration curve obtained through local regression of the observed pairwise treatment effect on the predicted pairwise treatment effect. The observed benefits are regressed on the predicted benefits using a locally weighted scatterplot smoother (loess). It is a non-parametric regression technique used to create a smoothing function by fitting a curve through a scatterplot of data points. It combines multiple local regressions to fit a curve to subsets of the data, allowing it to capture complex patterns without assuming a specific global model.
The Integrated Calibration-for-Benefit ($\text{ICI}_{\text{benefit}})$ (Rekkas et al., 2023; Maas et al., 2023) measures the average difference between predicted and smooth observed benefit across different levels of prediction determined by the empirical density function of the predicted pairwise treatment effect. $\text{E50}_{\text{benefit}}$ represents the $50^{th}$ percentile of the absolute distance between predicted and smooth observed benefit. It assesses how well the predicted treatment effect matches the observed treatment effect at the median. It offers insight into the model’s calibration at the central point rather than across the entire prediction range. For these measures, null values indicate perfect calibration, meaning that the model’s predicted treatment benefits align perfectly with the observed benefits.
@@ 3.5.4 Root mean squared error
with $\theta_{i}$ the true treatment benefit for individual $i$ at time $t$. For simulated data, $\hat{\theta}_{i}$ is the predicted benefit for individual $i$, that is, the difference in survival probabilities at a pre-specified time point $t$ for a patient with and without treatment given the covariate values:
@@ 4.1.1 Data generation process 1
In Table 1 and Figure 5, the highest average $\text{C}_{\text{benefit}}$ were obtained with the IF model. It was also the closest to the oracle value. ALASSO obtained the lowest value in terms of $\text{C}_{\text{benefit}}$, which could indicate that the linear ALASSO model has difficulty capturing the treatment by biomarker interactions, whereas IF, due to its specific design for such interactions, demonstrates superior performance in modeling them.
$\text{E50}_{\text{benefit}}$ and $\text{ICI}_{\text{benefit}}$ values remained substantially above zero across all models, which could highlight difficulties in obtaining solid calibration or limitations of the matching procedure between control group patients and treatment group patients. In contrast, RMSE values were not too far from zero, particularly for ALASSO. Note that for ALASSO, a small RMSE may not be so clinically useful if the C-for-benefit is close to 0.5.
If we only consider the machine learning methods, IF performed well in terms of discrimination performances and accuracy (higher concordance indices and lower RMSE), while FNNs had better calibration performances (lower $\text{E50}_{\text{benefit}}$ and $\text{ICI}_{\text{benefit}}$). Notably, CoxCC and CoxTime yielded similar results, as expected, since no interaction with time was generated in the simulation setup.
Table 1 also presents the results of models trained on an expanded cohort of patients ($n=10,000$). The results remain largely consistent between $n=1,000$ and $n=10,000$. For the machine learning methods, both the C-statistics and RMSE show only minor improvements compared to models trained on a smaller cohort. Meanwhile, $\text{C}_{\text{benefit}}$ decreases slightly, whereas $\text{E50}_{\text{benefit}}$ and $\text{ICI}_{\text{benefit}}$ exhibit slight increases.
@@
Model C $\text{C}_{\text{benefit}}$ $\text{E50}_{\text{benefit}}$ $\text{ICI}_{\text{benefit}}$ RMSE C $\text{C}_{\text{benefit}}$ $\text{E50}_{\text{benefit}}$ $\text{ICI}_{\text{benefit}}$ RMSE
@@
We also compared the models at different horizon points (Supplementary Material Table 7) and for different levels of censoring (Supplementary Material Table 8, Setting 1). All the results were improved when the censoring rate was lower ($20\%$ instead of $50\%$). Regarding the choice of the horizon time, the oracle values decreased over time, while the models performed similarly between the first and third quartiles of the time range.
The models were also compared using a matching procedure based on patients’ characteristics (Table 8, Setting 2). All the models performed worse compared to using a matching procedure based on predicted treatment benefit. As suggested by Klaveren et al. (2017), this is expected as the predicted treatment benefit for $2$ patients within a pair matched by patient characteristics tends to be less similar.
@@ 4.1.2 Data generation process 2
In Table 2, CoxCC, CoxTime, and IF exhibit comparable performances, with C values close to each other. CoxTime achieves the highest $\text{C}_{\text{benefit}}$ at $0.616$, with CoxCC and IF closely following. CoxTime also yields the lowest $\text{E50}_{\text{benefit}}$ and $\text{ICI}_{\text{benefit}}$, at $0.134$ and $0.153$. ALASSO consistently shows less good performance metrics compared to the other models.
Table 2 also presents the results for the model trained on a larger sample size. The results demonstrate only subtle differences across the evaluated metrics compared to the results obtained with the smaller sample size.
@@
Model C $\text{C}_{\text{benefit}}$ $\text{E50}_{\text{benefit}}$ $\text{ICI}_{\text{benefit}}$ RMSE C $\text{C}_{\text{benefit}}$ $\text{E50}_{\text{benefit}}$ $\text{ICI}_{\text{benefit}}$ RMSE
@@
Figure 5: $\text{C}_{\text{benefit}}$ and RMSE values for each model in data generation process 2. Each boxplot displays $100$ points corresponding to each simulated set.
4.2 Results on the three cancer data sets
FNN-based methods, IF and ALASSO were applied to $3$ cancer clinical trial data datasets. For the first two datasets, which involved early breast cancer, all comparison measures were computed on the test set at horizons of 2 and 5 years, as these are standard time points of interest for clinical investigators. The third dataset consisted of renal carcinoma patients, with comparison measures assessed at 1 year, as the endpoint is progression-free survival (PFS) in the advanced setting. We applied bootstrap resampling $M=100$ times to each dataset and computed confidence intervals for all evaluation measures.
4.2.1 Conditional average treatment effects of taxane chemotherapy
The bootstrapped evaluation measures at 2 and 5 years are presented in Table 3.
Overall, the neural network–based models adapted for survival, CoxCC and CoxTime, achieved the highest concordance values at 2-years, indicating superior discriminative ability compared to ALASSO and IF. When focusing on benefit-specific measures, IF slightly outperformed other models in terms of $\text{C}_{\text{benefit}}$, $\text{ICI}_{\text{benefit}}$, and $\text{E50}_{\text{benefit}}$, at both 2 and 5 years, suggesting it more accurately captures treatment effect heterogeneity. These results suggest that while neural networks are effective at predicting overall survival in this case study, they may not fully capture treatment interactions. In contrast, IF, designed to model individualized treatment effects, better identifies who benefits most from taxane chemotherapy.
Table 3: Bootstrapped evaluation measures at 2 and 5 years for the meta-analysis in the effect of taxane chemotherapy. - matching on predicted benefit.
@@ 4.2.1 Conditional average treatment effects of taxane chemotherapy
Model C $\text{C}_{\text{benefit}}$ $\text{E50}_{\text{benefit}}$ $\text{ICI}_{\text{benefit}}$ C $\text{C}_{\text{benefit}}$ $\text{E50}_{\text{benefit}}$ $\text{ICI}_{\text{benefit}}$
ALASSO $\underset{[0.499,0.563]}{0.511}$ $\underset{[0.514,0.750]}{0.650}$ $\underset{[0.038,0.212]}{0.127}$ $\underset{[0.063,0.206]}{0.138}$ $\underset{[0.15,0.905]}{0.563}$ $\underset{[0.491,0.758]}{0.632}$ $\underset{[0.032,0.19]}{\mathbf{0.099}}$ $\underset{[0.062,0.189]}{0.117}$
CoxCC $\underset{[0.632,0.714]}{\mathbf{0.673}}$ $\underset{[0.515,0.743]}{0.633}$ $\underset{[0.047,0.29]}{0.164}$ $\underset{[0.089,0.265]}{0.173}$ $\underset{[0.214,0.809]}{0.552}$ $\underset{[0.456,0.711]}{0.591}$ $\underset{[0.04,0.23]}{0.119}$ $\underset{[0.068,0.21]}{0.132}$
CoxTime $\underset{[0.629,0.713]}{\mathbf{0.673}}$ $\underset{[0.482,0.753]}{0.616}$ $\underset{[0.059,0.297]}{0.168}$ $\underset{[0.09,0.273]}{0.174}$ $\underset{[0.25,0.787]}{\mathbf{0.565}}$ $\underset{[0.460,0.692]}{0.577}$ $\underset{[0.045,0.23]}{0.123}$ $\underset{[0.062,0.218]}{0.131}$
IF $\underset{[0.454,0.577]}{0.506}$ $\underset{[0.554,0.760]}{\mathbf{0.656}}$ $\underset{[0.029,0.203]}{\mathbf{0.103}}$ $\underset{[0.047,0.19]}{\mathbf{0.109}}$ $\underset{[0.127,0.891]}{0.505}$ $\underset{[0.551,0.761]}{\mathbf{0.652}}$ $\underset{[0.033,0.174]}{\mathbf{0.099}}$ $\underset{[0.056,0.181]}{\mathbf{0.106}}$
@@
4.2.2 Conditional average treatment effects of adjuvant trastuzumab
For the second breast cancer data set, we observed in Table 4 that CoxCC and CoxTime reached the highest concordance values. Regarding treatment benefit, the machine-learning-based methods obtained higher values than ALASSO. CoxCC and CoxTime achieved the lowest $\text{E50}_{\text{benefit}}$ and $\text{ICI}_{\text{benefit}}$, indicating superior calibration of predicted benefit compared to ALASSO and IF. On the other hand, IF provided the highest $\text{C}_{\text{benefit}}$, suggesting it captures treatment heterogeneity more effectively, although at the cost of reduced calibration.
Table 4: Bootstrapped evaluation measures at 2 and 5 years in the RCT on adjuvant trastuzumab. - Matching on predicted benefit.